Grade 12Mechanics

A hollow smooth uniforms sphere A of mass M tools without sliding on a smooth horizontal surface . It collides head on elastically with another stationary smooth solid sphere B of mass M and same radius. What will be the ratio of K.E. of B to that of A just after collision??

supreet dey
8 Years agoGrade 12

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
when two identical bodies collide & one body is at rest initially then 

linear momentam is totally transfered to other body...

let the velocity of sphere A initially was V....

final velocity of A= 0

final velocity of other sphere B is V ....   (linear momentam is transfered)

kinetic energy of B =mV2/2 ..............1

kinetic energy of A = IW2/2

=mV2/3 ................2    (Isphere = 2/3 mr2 )

eq1 / eq2 =3/2

 

Regards

Arun (askIITians forum expert)